# On dense subalgebras of the singular ideal in groupoid C*-algebras

### v1

@@ Abstract.
We prove that ideals in amenable second-countable non-Hausdorff étale groupoid $C^{*}$-algebras are determined by their isotropy fibres. As an application, we characterise when the singular functions in Connes’ algebra are dense in the singular ideal in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
@@ Introduction
In recent years non-Hausdorff groupoids and their $C^{*}$-algebras have gained increased attention. Many groupoids arising from dynamics and geometry are non-Hausdorff, and it is therefore important to develop a robust theory. Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations which are non-Hausdorff ([10], [21], [22], [23]).
In contrast with the Hausdorff case, the functions in the reduced $C^{*}$-algebra $C^{*}_{r}(G)$ of a non-Hausdorff groupoid $G$ are not necessarily continuous. Indeed, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$. The set of all such functions is known as the singular ideal $J$ in $C^{*}_{r}(G)$. Historically, the singular ideal has been an obstruction to understanding simplicity for reduced $C^{*}$-algebras of non-Hausdorff groupoids ([9]). Characterisations of simplicity and the ideal intersection property have been obtained for the quotient by this ideal ([9], [18]), which is known as the essential groupoid $C^{*}$-algebra ([19], [13]). Therefore, the only obstacle to understanding these properties for the reduced $C^{*}$-algebra is the singular ideal.
Much work has been done to find conditions that ensure the singular ideal $J$ vanishes (see [9], [19], [24], [15]). Recently in [16] vanishing of the singular ideal was characterised in terms of a property of certain collections of subgroups in isotropy groups of $G$. In [4] it is characterised when the singular ideal $J$ has trivial intersection with the underlying groupoid -algebra, $\mathscr{C}_{c}(G)$ known as Connes’ algebra (first defined in [10] for groupoids arising from foliations).
When the singular ideal is non-zero, less is known about its structure. For example, it is not known in general whether $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}$. This was shown for groupoids with finite “non-Hausdorfness” in [4] and in [16] the problem was shown to be equivalent to whether group $C^{*}$-algebras satisfy the Intersection Property. This reduction of the problem led to its solution for groupoids with isotropy groups that are direct limits of virtually torsion free solvable groups [16].
A related question asks for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$ ([4, Question 4.11(III)]). Progress could be made towards understanding the structure of $J$ if it has a dense subalgebra of functions in $\mathscr{C}_{c}(G)$ (since functions in $\mathscr{C}_{c}(G)$ are easier to understand). In recent work ([20]), it is shown that there exist non-amenable groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. The following question remains open however.
Question.
@@
The goal of our paper is to study the question highlighted above. Let us outline the main achievements of the paper. All results listed here are for amenable second-countable étale groupoids.
  • We show that ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. This generalises [8, Theorem 2.10] to the setting of non-Hausdorff groupoids.
  • Following the approach in [16], we characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras.
  • We describe some large classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. In particular, we show density holds for groupoids with abelian isotropy groups, and groupoids arising from contracting self-similar groups.
We now state our main results and describe them in more detail. Our first result concerns ideals in groupoid $C^{*}$-algebras. In [8] it is shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$. These images are known as the isotropy fibres of the ideal, and are denoted $(I_{x})_{x\in G^{(0)}}$. It is shown in [8, Theorem 2.10] that for amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres. We generalise this result to non-Hausdorff groupoids.
@@
We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid. The following Theorem reduces the density question for $J\cap\mathscr{C}_{c}(G)$ in $J$ to one about density of ideals in isotropy group $C^{*}$-algebras.
Theorem B (See Proposition 2.6).
The isotropy fibres $J_{x}$ are calculated explicitly in [16, Theorem 5.5]. Moreover, when $G$ is covered by countably many bisections and $G^{x}_{x}$ is amenable, it is shown $J_{x}$ equals the ideal $J_{G^{x}_{x},\mathcal{X}(x)}$ defined as the intersection of the kernels of quasi-regular representations associated to subgroups of $G^{x}_{x}$ in a certain a collection $\mathcal{X}(x)$. We refer the reader to Subsection 2.1 for the definition of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$. The following is then a rephrasing of Theorem B.
Theorem C (See Theorem 2.9).
@@
In the above theorem, the Density Property is a property of discrete amenable groups we introduce in Definition 2.8. Finite groups satisfy the Density Property automatically, and in Proposition 3.11 we show that all abelian groups have the Density Property. Currently, we do not know of any discrete amenable group that fails the Density Property. We remark that it is a corollary of Theorem C and a non-Hausdorff groupoid construction in [16, Section 6] that if there is such a group, then there is an amenable groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 2.12).
By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem C asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups. In light of this, we are able to describe some explicit classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem D (See Corollary 3.12).
Let $G$ be an amenable and second-countable étale groupoid. Assume for every $x\in G^{(0)}$ one of the following holds.
@@
    The isotropy group $G_{x}^{x}$ is abelian.
@@
    The subgroups $X\in\mathcal{X}(x)$ are finite.
@@
Theorem D covers many important classes of groupoids. In particular, any groupoid of germs associated with the action of a countable abelian group will satisfy (I). All groupoids with finite isotropy groups will satisfy (II). More generally, condition (II) is satisfied whenever the unit space of $G$ has finite source and range fibres.
In the setting of non-Hausdorff ample groupoids, one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [26]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$. A characterisation for the vanishing of $J_{\mathbb{C}}$ is given in [4], and together with [9] and [27] this leads to a complete characterisation of simplicity for Steinberg algebras (see also [28] and [15]). In this setting, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$? Indeed, this is presented as an open problem in [15]. In order to answer this question, we show it suffices to study density of $J\cap\mathscr{C}_{c}(G)$ in $J$.
@@
An important class of ample groupoids are those arising from self-similar groups (see [22]). These groupoids can be non-Hausdorff, and can also have non-vanishing singular ideal. Using our results, we answer the density question for groupoids arising from contracting self-similar groups.
Corollary F (See Corollary 4.5).
Let $G$ be the groupoid arising from the action of a contracting self-similar group on a finite alphabet. Then $J_{\mathbb{C}}$ is dense in $J$.
@@ Acknowledgments
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Topological groupoids and their $C^{*}$-algebras, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1. Hume was supported by NSERC Discovery Grant RGPIN-2021-03834. Gonzales would like to thank Xin Li for helpful discussions on ideas related to this paper.
@@ 1. Preliminaries
A topological groupoid $G$ is a groupoid equipped with a topology such that the product and inverse maps are continuous. By an étale groupoid $G$, we mean a locally compact topological groupoid for which the unit space $G^{(0)}$ is Hausdorff, and the range $r\colon G\to G^{(0)}$ and source $s\colon G\to G^{(0)}$ maps are local homeomorphisms. This paper will only consider étale groupoids, and primarily concerns non-Hausdorff groupoids (all results also hold for Hausdorff groupoids, but often for trivial reasons i.e. because the singular ideal vanishes for Hausdorff groupoids).
An open subset $U\subseteq G$ will be called an open bisection if the restrictions $r|_{U}\colon U\to r(U)$ and $s|_{U}\colon U\to s(U)$ are homeomorphisms. Note that any étale groupoid has a basis of open bisections. We define
@@
If the groupoid $G$ is Hausdorff, then $\mathscr{C}_{c}(G)=C_{c}(G)$, the set of all compactly supported continuous functions on $G$. However if $G$ is not Hausdorff, then functions in $\mathscr{C}_{c}(G)$ need not be continuous. We equip $\mathscr{C}_{c}(G)$ with the structure of a -algebra as follows. For $f,f_{1},f_{2}\in\mathscr{C}_{c}(G)$, and $g\in G$ define
@@
Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [6, Lemma 3.27], [25, Proposition II.4.2]). From now on, we will always treat elements of $C^{*}_{r}(G)$ as functions on $G$ in this way. Given $a\in C^{*}_{r}(G)$, we define
@@
The singular ideal $J$ is a (closed) ideal in $C^{*}_{r}(G)$. For $a\in C^{*}_{r}(G)$, ${\rm supp}^{\circ}(a)$ has empty interior in $G$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [1, Proposition 4.6], [4, Lemma 4.1(iii)]).

### v2

@@ Abstract.
We prove that ideals in amenable second-countable non-Hausdorff étale groupoid $C^{*}$-algebras are determined by their isotropy fibres. As an application, we characterise when the singular functions in Connes’ algebra are dense in the singular ideal in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras. We then show this density property holds for all $C^{*}$-algebras of groupoids with finite-by-nilpotent isotropy groups.
@@ Introduction
In recent years non-Hausdorff groupoids and their $C^{*}$-algebras have gained increased attention. Many groupoids arising from dynamics and geometry are non-Hausdorff, and it is therefore important to develop a robust theory. Indeed, there are natural examples of groupoids of germs, groupoids arising from self-similar groups, and groupoids arising from foliations that are non-Hausdorff ([11], [27], [28], [29]).
In contrast with the Hausdorff case, the functions in the reduced $C^{*}$-algebra $C^{*}_{r}(G)$ of a non-Hausdorff groupoid $G$ are not necessarily continuous. In fact, there can exist (non-zero) functions in $C^{*}_{r}(G)$ whose set of non-zero values has empty interior in $G$. The set of all such functions is known as the singular ideal $J$ in $C^{*}_{r}(G)$. Historically, the singular ideal has been an obstruction to understanding simplicity for reduced $C^{*}$-algebras of non-Hausdorff groupoids ([10]). Characterisations of simplicity and the ideal intersection property have been obtained for the quotient by this ideal ([10], [21]), which is known as the essential groupoid $C^{*}$-algebra ([22], [13]). Therefore, the singular ideal is the only obstacle to understanding these properties for $C^{*}_{r}(G)$.
Much work has been done to find conditions that ensure that the singular ideal $J$ vanishes (see [10], [22], [30], [15]). In [5, Theorem A] (see also [19, Theorem F]) it is characterised when the singular ideal has trivial intersection with the underlying groupoid -algebra $\mathscr{C}_{c}(G)$, known as Connes’ algebra (first defined in [11] for groupoids arising from foliations). Recently, [19, Theorem B] characterised vanishing of the singular ideal in terms of vanishing of explicit ideals in the isotropy group $C^{*}$-algebras.
When the singular ideal is non-zero, little is known about its structure. For example, it is not known in general whether $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}$. This implication was shown to hold for groupoids that are “finitely non-Hausdorff” in [5, Theorem C]. In [19], the problem was reduced to one about group $C^{*}$-algebras, and this led to a positive answer in many cases ([19, Theorem H]).
In order to understand the structure of $J$, it is necessary to ask the following deeper question (see [5, Question 4.11(III)]): for which groupoids is $J\cap\mathscr{C}_{c}(G)$ dense in $J$? In recent work ([25]), it was shown that there exist non-amenable étale groupoids for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. The following remains open however.
Question 1.
@@
A positive answer to this question would allow progress to be made towards understanding the structure of $J$ (since the functions in $\mathscr{C}_{c}(G)$ are easier to understand). It is therefore natural to ask a follow up question: can we explicitly describe elements of $J\cap\mathscr{C}_{c}(G)$?
The goal of our paper is to study these questions following the approach in [19] of studying the singular ideal through its isotropy fibres. Let us outline the main achievements and methods of the paper. The first four points listed here are for amenable and second-countable étale groupoids.
  • We show that ideals in a (non-Hausdorff) groupoid $C^{*}$-algebra are determined by their isotropy fibres. We derive this result from the Hausdorff case (proven in [9]) through a novel application of the Hausdorff cover (see [5]).
  • We use this result to characterise when $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ in terms of a property of explicit ideals in the isotropy group $C^{*}$-algebras. This reduces Question 1 to a property of discrete amenable groups, which we call the Density Property.
  • We show the Density Property is preserved under various operations on the group. We pair this with structural properties of the space of subgroups (equipped with the Chabauty topology) to prove that the Density Property holds for all finite-by-nilpotent groups.
  • This reduction allows us to prove that $J\cap\mathscr{C}_{c}(G)$ is dense in $J$ for groupoids with finite-by-nilpotent isotropy groups, groupoids arising from contracting self-similar group actions, and for amenable étale group bundles.
  • We describe an explicit family of functions whose span is always dense in $J\cap\mathscr{C}_{c}(G)$.
  • For ample groupoids, we completely describe the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$, and prove $J_{\mathbb{C}}$ is dense in $J\cap\mathscr{C}_{c}(G)$.
$\bullet$
Let us state our main results in more detail. The first four results describe the progress made towards answering Question 1. In [9] it was shown that ideals $I\trianglelefteq C^{*}(G)$ in a groupoid $C^{*}$-algebra are mapped to ideals in the isotropy group $C^{*}$-algebras by the restriction maps $C^{*}(G)\to C^{*}(G^{x}_{x})$. These images are known as the isotropy fibres of the ideal, and are denoted $(I_{x})_{x\in G^{(0)}}$. For amenable second-countable Hausdorff étale groupoids, ideals in the groupoid $C^{*}$-algebra are determined by their isotropy fibres (see [9, Theorem 2.10]). We generalise this result to non-Hausdorff groupoids.
@@
We apply this result to the singular ideal $J$ in the reduced $C^{*}$-algebra of a non-Hausdorff groupoid (see Proposition 3.1). The isotropy fibres $J_{x}$ are calculated explicitly in [19, Theorem 5.5], where it is shown that $J_{x}$ is equal to an intersection of kernels of quasi-regular representations associated with subgroups of $G^{x}_{x}$ in a certain collection $\mathcal{X}(x)$. This ideal is denoted $J_{G^{x}_{x},\mathcal{X}(x)}\trianglelefteq C^{*}(G^{x}_{x})$. We refer the reader to Definition 1.3 and Section 3 for the definitions of $\mathcal{X}(x)$ and $J_{G^{x}_{x},\mathcal{X}(x)}$.
Theorem B (See Theorem 3.4).
@@
In Theorem B, the Density Property is a property of discrete amenable groups introduced in Definition 3.3. Finite groups trivially satisfy the Density Property, and in Theorem 4.16 we show that all finite-by-nilpotent groups have the Density Property. Currently, we do not know of any discrete amenable group that fails to have the Density Property. We remark that it is a consequence of Theorem B and a non-Hausdorff groupoid construction in [19, Section 6] that if there exists such a group, then there also exists an amenable and second-countable étale groupoid for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$ (see Corollary 3.7).
By the definition of the ideals $J_{G^{x}_{x},\mathcal{X}(x)}$, Theorem B asserts that density of $J\cap\mathscr{C}_{c}(G)$ in $J$ can be understood entirely in terms of quasi-regular representations of isotropy groups. In light of this, we are able to describe some explicit classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem C (See Corollary 4.17).
Let $G$ be an amenable and second-countable étale groupoid. For each $x\in G^{(0)}$ let $\mathcal{X}(x)$ be as in Definition 1.3, and assume that one of the following holds.
@@
    The isotropy group $G_{x}^{x}$ is finite-by-nilpotent.
@@
    The subgroups $X\in\mathcal{X}(x)$ are all finite.
  3. (III)
    The subgroups $X\in\mathcal{X}(x)$ are all co-finite in $G_{x}^{x}$.
  4. (IV)
    The subgroups $X\in\mathcal{X}(x)$ are all normal in $G^{x}_{x}$.
@@
Here, a countable group $\Gamma$ is finite-by-nilpotent if there exists a finite normal subgroup $N\trianglelefteq\Gamma$ for which the quotient $\Gamma/N$ is nilpotent. Theorem C covers many important classes of groupoids. In particular, any groupoid of germs associated with the action of a countable finite-by-nilpotent group will satisfy (I) for all $x\in G^{(0)}$. Condition (II) is satisfied for all $x\in G^{(0)}$ whenever the closure of the unit space of $G$ has finite source and range fibres. This is the hypothesis assumed in [5, Theorem C] when proving the weaker property that $J\neq\{0\}$ implies $J\cap\mathscr{C}_{c}(G)\neq\{0\}.$
The next result considers étale group bundles – that is, étale groupoids whose range and source maps are equal. Historically, étale group bundles have proven to be a good resource for providing counterexamples to conjectures (see [18], [37] for example). It was shown in [25] that there exist non-amenable étale group bundles for which $J\cap\mathscr{C}_{c}(G)$ is not dense in $J$. A direct corollary of Theorem C (IV) is that density always holds for amenable étale group bundles.
Corollary D (See Corollary 4.18).
If $G$ is an amenable and second-countable étale group bundle, then $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
$\bullet$
The remaining results describe the structure of elements in $J\cap\mathscr{C}_{c}(G)$. In Section 5 we describe an explicit set whose span is dense in the subalgebra $J\cap\mathscr{C}_{c}(G)$. Elements in this spanning set are weighted sums of functions $\psi\in C_{c}\big(G^{(0)}\big)$ lifted by the source map, and are built using purely topological and algebraic data coming from the groupoid and its isotropy groups. This provides an explicit dense subspace of the singular ideal $J$ for groupoids as described in Theorem C.
@@
Let $G$ be an étale groupoid that is covered by countably many open bisections. For $x\in G^{(0)}$ let $b\colon G^{x}_{x}\to\mathbb{C}$ be a function with finite support $g_{1},\dots,g_{n}$ such that
(0.1) $\sum_{g\in kX}b(g)=0\text{ for all }k\in G^{x}_{x}\text{ and }X\in\mathcal{X}(x).$
Then, given any $h\in G_{x}$ and any open bisections $V_{1},\dots,V_{n}$ containing $hg_{1},\dots,hg_{n}$, there exists an (explicitly defined) open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that
(0.2) $f^{\psi}\coloneq\sum_{i=1}^{n}b(g_{i})(\psi\circ s|_{V_{i}})$
belongs to $J\cap\mathscr{C}_{c}(G)$ for all $\psi\in C_{c}(W)$. Furthermore, the linear span of such functions $f^{\psi}$ is dense in $J\cap\mathscr{C}_{c}(G)$ with respect to any $C^{*}$-norm.
We remark that an element $b\in\mathbb{C}[G^{x}_{x}]$ satisfies (0.1) if and only if, for every subgroup $X\in\mathcal{X}(x)$, $b$ belongs to the kernel of the associated quasi-regular representation $\lambda_{G^{x}_{x}/X}$.
$\bullet$
We now turn our attention to non-Hausdorff ample groupoids, where one can intersect $J$ with the complex Steinberg algebra $\mathbb{C}G$ (see [33]) to obtain what is known as the algebraic singular ideal $J_{\mathbb{C}}$. A characterisation for the vanishing of $J_{\mathbb{C}}$ is given in [5], and together with [10] and [34] this leads to a complete characterisation of simplicity for Steinberg algebras (see also [35] and [15]). In this paper we look beyond the vanishing of $J_{\mathbb{C}}$ and describe an explicit spanning set analogous to that in Theorem E. We stress the following: the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$ is now fully understood.
Theorem F (See Theorem 6.1).
Let $G$ be an ample groupoid that is covered by countably many open bisections. Let $x,b,h,V_{1},\dots,V_{n}$ be as in Theorem E. There exists a compact open neighbourhood $W\subseteq\bigcap_{i=1}^{n}s(V_{i})$ of $x$ such that the function
(0.3) $f^{W}\coloneq\sum_{i=1}^{n}b(g_{i})\mathbbm{1}_{U_{i}}$
belongs to $J_{\mathbb{C}}$, where $U_{i}\coloneq V_{i}\cap s^{-1}(W)$. Furthermore, $J_{\mathbb{C}}$ is equal to the linear span of functions of the form $f^{W}$.
For ample groupoids, one can ask the following density question: for which ample groupoids is $J_{\mathbb{C}}$ dense in $J$? Indeed, this is presented as an open problem in [15]. The following theorem implies that $J_{\mathbb{C}}$ is dense in $J$ if and only if $J\cap\mathscr{C}_{c}(G)$ is dense in $J$.
Theorem G (See Theorem 6.3).
@@
An important class of ample groupoids are those arising from self-similar groups (see [28]). These groupoids can be non-Hausdorff, and can also have non-vanishing singular ideal. For a groupoid arising from a contracting self-similar group, it was shown in [15] that $J_{\mathbb{C}}=\{0\}$ if and only if $J=\{0\}$. Using our results, we show that $J_{\mathbb{C}}$ is in fact dense in $J$ for all groupoids in this class.
Corollary H (See Corollary 6.4).
Let $G$ be the groupoid arising from a contracting self-similar group action. Then $J_{\mathbb{C}}$ is dense in $J$.
@@
$\bullet$
The paper is structured as follows. In Section 1, we give the necessary background on the Hausdorff cover groupoid and on isotropy fibres of ideals. In Section 2 we prove Theorem A, and then apply this result to the singular ideal $J$ in Section 3. Section 4 studies the Density Property for groups and ends by describing some classes of groupoids for which $J\cap\mathscr{C}_{c}(G)$ is dense in $J$. Sections 5 studies the structure of functions in $J\cap\mathscr{C}_{c}(G)$. Section 6 is exclusively concerned with ample groupoids and describes the structure of functions in the algebraic singular ideal $J_{\mathbb{C}}$.
@@ Acknowledgments
The authors would like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Topological groupoids and their $C^{*}$-algebras, where work on this paper was undertaken. This work was supported by EPSRC grant EP/V521929/1. Hume was supported by NSERC Discovery Grant RGPIN-2021-03834. Gonzales was supported by EPSRC grant EP/T517896/1. Gonzales would like to thank Xin Li for helpful discussions on topics related to this paper.
@@ 1. Preliminaries
A topological groupoid $G$ is a groupoid equipped with a topology such that the product and inverse maps are continuous. By an étale groupoid $G$, we mean a locally compact topological groupoid for which the unit space $G^{(0)}$ is Hausdorff, and the range $r\colon G\to G^{(0)}$ and source $s\colon G\to G^{(0)}$ maps are local homeomorphisms. Note that this ensures that the unit space $G^{(0)}$ is open in $G$. This paper will only consider étale groupoids, and primarily concerns non-Hausdorff groupoids (all results also hold for Hausdorff groupoids, but often for trivial reasons i.e. because the singular ideal vanishes for Hausdorff groupoids).
An open subset $U\subseteq G$ will be called an open bisection if the restrictions $r|_{U}\colon U\to r(U)$ and $s|_{U}\colon U\to s(U)$ are homeomorphisms. Note that the open bisections form a basis for the topology of any étale groupoid. We define Connes’ algebra as follows:
@@
If the groupoid $G$ is Hausdorff, then $\mathscr{C}_{c}(G)=C_{c}(G)$, the set of all compactly supported continuous functions on $G$. If $G$ is not Hausdorff, then the functions in $\mathscr{C}_{c}(G)$ need not be continuous. We equip $\mathscr{C}_{c}(G)$ with the structure of a -algebra as follows. For $f,f_{1},f_{2}\in\mathscr{C}_{c}(G)$, and $g\in G$ define
@@
Elements of the reduced groupoid $C^{*}$-alegbra $C^{*}_{r}(G)$ can be viewed as bounded Borel functions on $G$ in a canonical way (see [7, Lemma 3.15], [31, Proposition II.4.2]). From now on, we will always treat elements of $C^{*}_{r}(G)$ as functions on $G$ in this way. Given $a\in C^{*}_{r}(G)$, we define
@@
The singular ideal $J$ is a (closed) ideal in $C^{*}_{r}(G)$. For $a\in C^{*}_{r}(G)$, we have $a\in J$ if and only if $s\big({\rm supp}^{\circ}(a)\big)$ has empty interior in $G^{(0)}$ (see [2, Proposition 4.6], [5, Lemma 4.1(iii)]).
Now assume that the groupoid $G$ is covered by countably many open bisections. We say a unit $x\in G^{(0)}$ is Hausdorff if for any $g\in G_{x}^{x}\setminus\{x\}$ there exists an open neighbourhood $U$ of $g$ such that $U\cap G^{(0)}=\emptyset$. It is clear that $G$ is Hausdorff if and only if all units are Hausdorff. We let $C\subseteq G^{(0)}$ denote the set of Hausdorff units. By [22, Lemma 7.15], $C$ is dense in $G^{(0)}$ (the Hausdorff points complement the so-called dangerous points of [22]). An element $a\in C^{*}_{r}(G)$ belongs to the singular ideal $J$ if and only if $s\big({\rm supp}^{\circ}(a)\big)\subseteq G^{(0)}\setminus C$ (see [22, Proposition 7.18]).
$\bullet$